Operation Summary for March 15, 2010 - March 17, 2010 |
Monday March 15 The day shift began with Operations monitoring the Tevatron (TeV) store 7678 (79.0E30) and the Recycler stash (272.0E10), with the Antiproton Source stacking (9.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, repairs to the S:QP3 ‘coffin’ continued. Operations terminated store 7678 at 2:30 PM and then operators replaced the power supply for a TeV CAMAC crate. Operations put the TeV into HEP shot setup and established store 7679 at 3:53 PM with an Average Initial Luminosity (AIL) of 345.3E30. Preparations to reenergize QP3 were underway by the end of shift. During the evening shift, around 6:30 PM, the Linac H- Source output dropped off significantly. Linac personnel arrived, switched over to the I- Source, and tuned up. A TeV RF station (TRF1) tripped off on an incomplete gap envelope fault; operators turned off the heaters. Tuesday March 16 The midnight shift began with Operations monitoring store 7679 (111.07E30) and the stash (212.5E10), with Pbar stacking (23.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to the Meson MTest T932 experiment at 4 AM. Operations turned off the NuMI beam at 5 AM to prepare for an access later in the morning. Operations terminated store 7679 at 7:35 AM and put the TeV into shot setup. During the day shift, Operations turned off the Booster beam and prepared the area for access and for the 400 MeV area Debuncher water skid work. Operations established store 7680 at 9:03 AM with an AIL of 333.8E30. An MI access began at 9:30 AM. EE Support personnel replaced a regulation chassis for a MI quadrupole (Q847) power supply. Water group personnel found and repaired a LCW leak on a MI dipole (865-3) magnet. Booster, MI, and MiniBooNE personnel came out of access. Pbar resumed stacking and MiniBooNE resumed taking beam around 2:11 PM. Operations began sending beam to MTest at 2:29 PM. At 3:34 PM, the Booster Gradient Magnet Power Supply tripped off on a transformer indication; Booster personnel found a loose cable. During the evening shift, NuMI came out of access at 5 PM and was taking beam by 5:53 PM. TRF1 tripped off and caused a slight loss of luminosity. Operations disabled the MTest beam at 9 PM. Operations reported problems with a Pbar Debuncher RF station (DRF1-7). Wednesday March 17 The midnight shift began with Operations monitoring store 7680 (76.6E30) and the stash (225.2E10), with Pbar stacking (18.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:12 AM. At 6 AM, TeV personnel began conducting a crystal collimator study. Operations disabled the MTest beam at 7:15 AM for an access, but then a problem developed with the MT6 interlocks that inhibited beam. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow the following work: an F-Sector cryo system controls reboot; TRF1 work; MT6 interlock work; and an 8 GeV collimator study. |
|
More Information Comments and Suggestions |
| last modified 03/17/2010 email Fermilab |
| Security, Privacy, Legal | |